OpenStudy (anonymous):

Calculus question... Sketch the graph of a function that does not have an inflection point at a point where f''(x) = 0

OpenStudy (anonymous):

i was thinking, would f(x) = x^4 work? I mean it's second derivative would be a 12x^2 and that doesn't have an inflection point since it's a parabola.

OpenStudy (jamesj):

Yes, that works. Also a function where f''(x) = 0 everywhere, such as f(x) = x

OpenStudy (anonymous):

\[y=x^2\] would do

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

oh sorry. asked and answered. site is really really slow

OpenStudy (jamesj):

f(x) = x^2 is not an example as f''(x) = 2 everywhere.

OpenStudy (anonymous):

oh i get it! this makes sense now

OpenStudy (anonymous):

good point. sorry. \[f(x)=x\] \[f(x)=x^4\] etc